1-21 Understanding Tax Rate Concepts. Indicate whether the following statements are true r false and, if false, explain why a. b. c. Tax-exempt income would cause the taxpayers average tax rate to increase. Tax-exempt income would cause the taxpayers marginal tax rate to decrease Tax-exempt income would cause the taxpayers effective tax rate to decrease
Answer #1

a. Tax-exempt income would cause the taxpayer's average tax rate to increase - False

Reason: By earning tax-exempt income, we are increasing the denominator only in the equation to calculate average rate. The numerator, i.e, tax payable remains the same. Rather tax-exempt income would result in decrease in the average tax rate.

Example: Tax on Non exempt income = 50,000; Taxable Income = 200,000 Average rate of tax = 50,000 / 200,000 = 25%

If say the person now earns 30,000 of tax-exempt income, average rate of tax = (50,000 + 0) / (200,000+30,000) = 21.74%

b. Tax-exempt income would cause the taxpayer's marginal tax rate to decrease - False

Reason - Marginal tax rate would be the amount of tax payable on per 1 unit increase in income. Typically applicable in case of individuals, various slabs are applicable to individuals.  Tax exempt income is not considered to consider next slab of rate.

Example: Say Mr. X is in the tax slab of 0% up to 250,000, 10% up to 500,000. 20% up to 10,00,000 and 30% above that.

Now he is earning Rs 8,00,000 out of which 4,00,000 is the tax-exempt income. He'll be considered in the second slab and the tax-exempt income would have no bearing of selections of slabs

c. Tax-exempt income would cause the taxpayers's effective tax rate to decrease - False

Reason: The formula for calculating effective tax rate is: Tax Expenses / Taxable Income

Continuing with the example quoted in a. above, effective tax rate would be 25%.

Continuing with the example quoted in b. above, effective tax rate for Mr. X would be calculated as follows:

Tax expenses = (400,000 - 250,000) x 10% = 15,000; Taxable Income = 400,000

Effective tax rate = 15,000 / 400,000 = 3.75%